# Democratizing Tabular Data Access with an Open‑Source Synthetic‑Data SDK

### v1

@@ I Introduction
The development of Machine Learning applications requires broad access to training data. This necessity has become more critical in recent years with the advent of Deep Learning, which requires large-scale datasets to effectively train models. At the same time, the volume of data produced has grown rapidly, as companies and institutions now routinely collect vast amounts of data from user interactions and deployed applications. However, this exponential growth has not translated into equivalent accessibility. Privacy regulations, proprietary restrictions, and ethical considerations increasingly restrict the availability and usage of real-world datasets for development. This phenomenon, referred to as the data bottleneck [1], affects multiple domains and data modalities with tabular datasets being no exception.
In this context, synthetic data arises as a promising solution to fill the data availability gap, offering means to securely access high-quality datasets, guaranteeing their utility and minimizing the privacy risks associated with data sharing [2, 3, 4, 5]. Synthetic datasets emerge as a versatile tool, also providing solutions for numerous problems other than privacy. For instance, scaling data volume, addressing class imbalance, imputing missing data or enabling fairness-aware evaluations [6]. Tackling these problems as a whole dramatically increases the usability and practicality of synthetic data [7]. However, despite the growing body of literature leveraging Deep Learning for tabular data synthesis, many methods fall short of providing a holistic solution that addresses multiple aspects of data synthesis simultaneously, especially considering challenges like data heterogeneity, fidelity, and privacy protection [8, 9].
@@ I-A Contributions
  • High data quality with a simple design: Based on TabularARGN [10], the SDK delivers synthetic data quality on a par with state-of-the-art (SOTA) models without compromising simplicity. Additionally, it achieves substantial speed-ups in training and generation times over current benchmarks.
  • Privacy and fairness features: The framework considers privacy-preserving value ranges for sampling, has built-in regularization layers and early stopping for training, and provides training mechanisms with differencial privacy (DP) guarantees [11]. In generation, it supports the incorporation of fairness across sensitive attributes [12].
  • Automated quality assurance: The SDK provides a built-in quality assurance module [13], generating standardized, in-depth reports that quantitatively assess both privacy and data fidelity of synthetic datasets.
  • Open source package: Fully written in Python, the SDK can be locally installed via pip. The package is open-sourced under a fully permissive Apache Version 2.0 Open Source license111 https://github.com/mostly-ai/mostlyai.
@@ II Architecture Overview and Workflow
In this Section we present an overview of the SDK. Built to support diverse usage scenarios, the SDK enables users from diverse backgrounds, from academic researchers to data science practitioners, to programmatically generate, share, and manipulate high-quality synthetic datasets. The toolkit provides a clearly structured interface. It is freely available as an open-source package under the permissive Apache Version 2.0 license, and can be easily installed via pip install -U ’mostlyai[local]’ (requiring Python 3.10$\geq$ 3.10≥ 3.10).
@@ II-A Core Components of the SDK
The evaluation framework [13] uses a holdout-based approach to compare synthetic and original data, covering a range of quantitative metrics such as distributional similarity, embedding-based measures, and nearest-neighbor distances. It supports diverse data types, including sequential and contextual features.
@@ III-A Data Preparation
Before training, raw input data undergoes a comprehensive pre-processing pipeline that includes automatic type detection, encoding assignment, and privacy-preserving value protection mechanisms. The SDK follows the data encoding protocol proposed by [10], where all columns in the original dataset are transformed into one or more categorical sub-columns. As a result, the encoded dataset exclusively consists of categorical variables, regardless of the domain of the original dataset. Table I summarizes all the supported tabular encoding types.
@@ IV-A Auto-regressive Tabular Model Overview
The generator artifact is trained using the Tabular Auto-Regressive Generative Network (TabularARGN), detailed in [10]. Its architecture provides a shallow, any-order auto-regressive model that operates on discretized features and is optimized using categorical cross-entropy loss. At its core, TabularARGN maps the generation process to an auto-regressive framework, estimating the joint probability distribution of the data as a product of simpler conditional probabilities. To handle multi-table datasets, the model extends the auto-regressive conditional probability framework to capture dependencies across related tables defined by the data schema. In these scenarios, the SDK explicitly defines an ordering in which tables are processed, establishing parent-children relationships between them. This way, records from parent tables are compressed into context embeddings, which are then used for conditional generation of their children tables.
@@ IV-C Privacy Mechanisms
  • Automatic Epoch Limiting: The maximum number of training epochs is automatically restricted based on the number of available training samples. This ensures that the training procedure does not have the opportunity to memorize individual records.
@@ V-A Flexible Generation
Rebalancing enables users to address class imbalance by generating additional synthetic samples for minority groups in the data. By specifying desired class proportions, users can increase the representation of underrepresented outcomes while ensuring that the remaining features remain consistent and realistic. In comparative evaluations [14], TabularARGN-based upsampling has demonstrated advantages over traditional techniques such as naive oversampling [15] and SMOTE-NC [16]. Unlike these methods, which often duplicate existing records or interpolate between them, TabularARGN generates a diverse set of realistic synthetic samples that more accurately reflect the underlying data distribution, leading to better performance for machine learning model trained on these balanced datasets, particularly when the minority class is very small  [^2] See the SDK usage tutorial for instructions on rebalancing: https://github.com/mostly-ai/mostlyai/tree/main/docs/tutorials/rebalancing.
@@
This is achieved by sampling the target column last and adjusting its conditional probabilities based on the previously generated sensitive attributes, thus enforcing statistical parity. As a result, users can create fairness-aware datasets, allowing downstream models to produce fairer, more balanced predictions, even when evaluated on real, biased data. These constraints are applied dynamically during generation, making it possible to flexibly manage the trade-off between fairness and accuracy without retraining the generator [12] [^4] To learn more about using the fairness feature, check out the SDK tutorial at https://github.com/mostly-ai/mostlyai/tree/main/docs/tutorials/fairness..
@@
In addition to flexible generation order, the engine allows users to further control the sampling process by adjusting the temperature and top-p parameters. Lowering the temperature (e.g. <1<1< 1) makes the model more likely to generate common values, which is useful when users want output that strictly follows known rules or business logic, for example, generating only valid product codes or adhering closely to typical user behaviors. The rise in temperature (e.g. >1>1> 1) increases the variety of generated data, which can be valuable for stress testing models or exploring rare scenarios.
@@ VI Empirical Performance and Community Adoption
The SDK offers a practical solution for a variety of data challenges, including privacy-safe synthesis, rebalancing of skewed datasets, imputation of missing values, and fair synthetic data. To comprehensively assess the real-world impact of the SDK we elaborate on two complementary aspects of our work. On the one hand, the empirical evidence presented by [10] is summarized, highlighting the strong performance of TabularARGN, the core model of the SDK, in terms of generation accuracy, robust privacy guarantees, and efficient runtime. On the other hand, the growing community adoption observed since the SDK’s public release in January 2025 is discussed, demonstating its practicality, usability, and relevance.
@@ VI-A Synthetic Data Quality
To demonstrate the effectiveness of our approach, we summarize key empirical results previously reported in [10], where TabularARGN was benchmarked against state-of-the-art methods for both flat and sequential tabular datasets. Evaluations were carried out following the protocol proposed by [17], which assesses the quality of synthetic data in terms of both statistical fidelity, measured through low-order marginal statistics such as univariate and bivariate accuracies, and privacy protection quality, quantified via metrics like Distance to Closest Record (DCR) share. This evaluation protocol is implemented in an open-source package [13] available at [18]. The SDK leverages this package to produce the comprehensive quality assurance reports described previously in Section 2.
For flat tables, the model was compared with methods such as CTGAN [19], STaSy [20], TabSyn [21], TabMT [22], and Unmasking Trees [23]. For sequential data, benchmarks included REalTabFormer [24], RC-TGAN [25], and ClavaDDPM [26]. Although several datasets were examined, in this Section we only present results of one dataset for each data regime, the adult census dataset and the baseball dataset, comprising longitudinal player-performance statistics. Figure 4 shows the main results, which underscore TabularARGN’s ability to achieve state-of-the-art performance while providing significant speed-ups over the benchmarks, even when training with DP, which can negatively impact accuracy and runtime.
@@ VI-B Scalability and Complex Data Handling
Despite significant progress in the field, the majority of deep learning–based tabular data generators remain focused on solving only subsets of problems, such as privacy or fidelity, often at the expense of flexibility, scalability, or support for real-world database topologies. A prominent example is the Synthetic Data Vault (SDV) [27], a widely used Business Source Python library developed by DataCebo, which offers tools for single-table, multi-table, and time-series data synthesis, leveraging a range of generative modeling techniques including GANs and VAEs. However, the benchmarking and experimental assessments  [^5] SDV vs. SDK: Single table scenario at https://mostly.ai/blog/a-comparison-of-synthetic-data-vault-and-mostly-ai-part-1-single-table-scenario show that while SDV excels in certain single-table use cases, it can encounter scalability and performance limitations with larger or more complex datasets, and may require manual configuration or extensive parameter tuning. Extending this analysis to a two-table dataset with complex relational and temporal dependencies, we observed that SDV struggled to capture these patterns, whereas the SDK handled them more effectively, achieving higher fidelity and utility under these conditions  [^6] SDV vs. SDK: Sequential Scenario at https://mostly.ai/blog/a-comparison-of-synthetic-data-vault-and-mostly-ai-part-2-sequential-scenario .

### v2

# Democratizing Tabular Data Access with an Open-Source Synthetic-Data SDK

@@ I Introduction
The development of Machine Learning applications requires broad access to training data. This necessity has become more critical in recent years with the advent of Deep Learning, which requires large-scale datasets to effectively train models. At the same time, the volume of data produced has grown rapidly, as companies and institutions now routinely collect vast amounts of data from user interactions and deployed applications. However, this exponential growth has not translated into equivalent accessibility. Privacy regulations, proprietary restrictions, and ethical considerations increasingly restrict the availability and usage of real-world datasets for development. This phenomenon, referred to as the data bottleneck [Shani_2023], affects multiple domains and data modalities with tabular datasets being no exception.
In this context, synthetic data arises as a promising solution to fill the data availability gap, offering means to securely access high-quality datasets, guaranteeing their utility and minimizing the privacy risks associated with data sharing [drechsler2011synthetic, jordon2022synthetic, united2023synthetic, hu2023sokprivacypreservingdatasynthesis]. Synthetic datasets emerge as a versatile tool, also providing solutions for numerous problems other than privacy. For instance, scaling data volume, addressing class imbalance, imputing missing data or enabling fairness-aware evaluations [NEURIPS2021_ba9fab00]. Tackling these problems as a whole dramatically increases the usability and practicality of synthetic data [vanbreugel2023privacynavigatingopportunitieschallenges]. However, despite the growing body of literature leveraging Deep Learning for tabular data synthesis, many methods fall short of providing a holistic solution that addresses multiple aspects of data synthesis simultaneously, especially considering challenges like data heterogeneity, fidelity, and privacy protection [bauer2024, shi2025comprehensivesurveysynthetictabular].
@@ I-A Contributions
  • High data quality with a simple design: Based on TabularARGN [tabularargn], the SDK delivers synthetic data quality on a par with state-of-the-art (SOTA) models without compromising simplicity. Additionally, it achieves substantial speed-ups in training and generation times over current benchmarks.
  • Privacy and fairness features: The framework considers privacy-preserving value ranges for sampling, has built-in regularization layers and early stopping for training, and provides training mechanisms with differencial privacy (DP) guarantees [dwo14-dp]. In generation, it supports the incorporation of fairness across sensitive attributes [krc23].
  • Automated quality assurance: The SDK provides a built-in quality assurance module [qareport], generating standardized, in-depth reports that quantitatively assess both privacy and data fidelity of synthetic datasets.
  • Open source package: Fully written in Python, the SDK can be locally installed via pip. The package is open-sourced under a fully permissive Apache Version 2.0 Open Source license [^1] https://github.com/mostly-ai/mostlyai.
@@ II Architecture Overview and Workflow
In this Section we present an overview of the SDK. Built to support diverse usage scenarios, the SDK enables users from diverse backgrounds, from academic researchers to data science practitioners, to programmatically generate, share, and manipulate high-quality synthetic datasets. The toolkit provides a clearly structured interface. It is freely available as an open-source package under the permissive Apache Version 2.0 license, and can be easily installed via pip install -U ’mostlyai[local]’ (requiring Python $\geq 3.10$).
@@ II-A Core Components of the SDK
The evaluation framework [qareport] uses a holdout-based approach to compare synthetic and original data, covering a range of quantitative metrics such as distributional similarity, embedding-based measures, and nearest-neighbor distances. It supports diverse data types, including sequential and contextual features.
@@ III-A Data Preparation
Before training, raw input data undergoes a comprehensive pre-processing pipeline that includes automatic type detection, encoding assignment, and privacy-preserving value protection mechanisms. The SDK follows the data encoding protocol proposed by [tabularargn], where all columns in the original dataset are transformed into one or more categorical sub-columns. As a result, the encoded dataset exclusively consists of categorical variables, regardless of the domain of the original dataset. Table I summarizes all the supported tabular encoding types.
@@ IV-A Auto-regressive Tabular Model Overview
The generator artifact is trained using the Tabular Auto-Regressive Generative Network (TabularARGN), detailed in [tabularargn]. Its architecture provides a shallow, any-order auto-regressive model that operates on discretized features and is optimized using categorical cross-entropy loss. At its core, TabularARGN maps the generation process to an auto-regressive framework, estimating the joint probability distribution of the data as a product of simpler conditional probabilities. To handle multi-table datasets, the model extends the auto-regressive conditional probability framework to capture dependencies across related tables defined by the data schema. In these scenarios, the SDK explicitly defines an ordering in which tables are processed, establishing parent-children relationships between them. This way, records from parent tables are compressed into context embeddings, which are then used for conditional generation of their children tables.
@@ V-A Flexible Generation
Rebalancing enables users to address class imbalance by generating additional synthetic samples for minority groups in the data. By specifying desired class proportions, users can increase the representation of underrepresented outcomes while ensuring that the remaining features remain consistent and realistic. In comparative evaluations [upsampling], TabularARGN-based upsampling has demonstrated advantages over traditional techniques such as naive oversampling [imblearn_naive] and SMOTE-NC [imblearn_smotenc]. Unlike these methods, which often duplicate existing records or interpolate between them, TabularARGN generates a diverse set of realistic synthetic samples that more accurately reflect the underlying data distribution, leading to better performance for machine learning model trained on these balanced datasets, particularly when the minority class is very small  [^2] See the SDK usage tutorial for instructions on rebalancing: https://github.com/mostly-ai/mostlyai/tree/main/docs/tutorials/rebalancing.
@@
This is achieved by sampling the target column last and adjusting its conditional probabilities based on the previously generated sensitive attributes, thus enforcing statistical parity. As a result, users can create fairness-aware datasets, allowing downstream models to produce fairer, more balanced predictions, even when evaluated on real, biased data. These constraints are applied dynamically during generation, making it possible to flexibly manage the trade-off between fairness and accuracy without retraining the generator [krc23] [^4] To learn more about using the fairness feature, check out the SDK tutorial at https://github.com/mostly-ai/mostlyai/tree/main/docs/tutorials/fairness..
@@
In addition to flexible generation order, the engine allows users to further control the sampling process by adjusting the temperature and top-p parameters. Lowering the temperature (e.g. <1<1) makes the model more likely to generate common values, which is useful when users want output that strictly follows known rules or business logic, for example, generating only valid product codes or adhering closely to typical user behaviors. The rise in temperature (e.g. >1>1) increases the variety of generated data, which can be valuable for stress testing models or exploring rare scenarios.
@@ VI Empirical Performance and Community Adoption
The SDK offers a practical solution for a variety of data challenges, including privacy-safe synthesis, rebalancing of skewed datasets, imputation of missing values, and fair synthetic data. To comprehensively assess the real-world impact of the SDK we elaborate on two complementary aspects of our work. On the one hand, the empirical evidence presented by [tabularargn] is summarized, highlighting the strong performance of TabularARGN, the core model of the SDK, in terms of generation accuracy, robust privacy guarantees, and efficient runtime. On the other hand, the growing community adoption observed since the SDK’s public release in January 2025 is discussed, demonstating its practicality, usability, and relevance.
@@ VI-A Synthetic Data Quality
To demonstrate the effectiveness of our approach, we summarize key empirical results previously reported in [tabularargn], where TabularARGN was benchmarked against state-of-the-art methods for both flat and sequential tabular datasets. Evaluations were carried out following the protocol proposed by [pla21-qa], which assesses the quality of synthetic data in terms of both statistical fidelity, measured through low-order marginal statistics such as univariate and bivariate accuracies, and privacy protection quality, quantified via metrics like Distance to Closest Record (DCR) share. This evaluation protocol is implemented in an open-source package [qareport] available at [mostlyai-qa]. The SDK leverages this package to produce the comprehensive quality assurance reports described previously in Section 2.
For flat tables, the model was compared with methods such as CTGAN [xu19], STaSy [kim23-StaSy], TabSyn [zha24-TabSyn], TabMT [gul23-TabMT], and Unmasking Trees [mcc24-UnmaskingTrees]. For sequential data, benchmarks included REalTabFormer [sol23-Realtabformer], RC-TGAN [gue23-rctgan], and ClavaDDPM [pan24-clavaddpm]. Although several datasets were examined, in this Section we only present results of one dataset for each data regime, the adult census dataset and the baseball dataset, comprising longitudinal player-performance statistics. Figure 4 shows the main results, which underscore TabularARGN’s ability to achieve state-of-the-art performance while providing significant speed-ups over the benchmarks, even when training with DP, which can negatively impact accuracy and runtime.
@@ VI-B Scalability and Complex Data Handling
Despite significant progress in the field, the majority of deep learning–based tabular data generators remain focused on solving only subsets of problems, such as privacy or fidelity, often at the expense of flexibility, scalability, or support for real-world database topologies. A prominent example is the Synthetic Data Vault (SDV) [pat16-sdv], a widely used Business Source Python library developed by DataCebo, which offers tools for single-table, multi-table, and time-series data synthesis, leveraging a range of generative modeling techniques including GANs and VAEs. However, the benchmarking and experimental assessments  [^5] SDV vs. SDK: Single table scenario at https://mostly.ai/blog/a-comparison-of-synthetic-data-vault-and-mostly-ai-part-1-single-table-scenario show that while SDV excels in certain single-table use cases, it can encounter scalability and performance limitations with larger or more complex datasets, and may require manual configuration or extensive parameter tuning. Extending this analysis to a two-table dataset with complex relational and temporal dependencies, we observed that SDV struggled to capture these patterns, whereas the SDK handled them more effectively, achieving higher fidelity and utility under these conditions  [^6] SDV vs. SDK: Sequential Scenario at https://mostly.ai/blog/a-comparison-of-synthetic-data-vault-and-mostly-ai-part-2-sequential-scenario .